At what temperature is the Fahrenheit scale reading equal to (a) twice that of the Celsius and (b) half that of the Celsius?

Grade:upto college level

5 Answers

mazidur rahman
7 years ago
at -24.613 F temperature is the Fahrenheit scale reading equal to (a) twice that of the Celsius and at 160 F temperatuar is the Fahrenheit scale reading equal to (b) half that of the Celsius?
mazidur rahman
7 years ago
at -24.613 F temperature is the Fahrenheit scale reading equal to twice that of the Celsius and at 160 F temperatuar is the Fahrenheit scale reading equal to half that of the Celsius?
khadija
5 years ago

asic conversion formula: F = 9/5C + 32
Our problem: 2C = F

Substitute F in the first formula with 2C from the second, which gives you:
2C = 9/5C + 32

Subtract 9/5C from both sides, which gives:
2C - 9/5C = 32

Simplify the left side (2C = 10/5C): 10/5C - 9/5C = 32, which equals 1/5C = 32

Multiply 5 to each side, which leaves C = 160

So, 160 degrees Celsius will equal 320 degrees Fahrenheit.

 

khadija
5 years ago
 
At what temperature is the Fahrenheit scale reading equal to twice that of the Celsius?
 

C = 5/9(F-32) or (F-32)/1.8

We want C = 2F or F = C/2

So C = (C/2 - 32)/1.8

1.8 C = C/2 - 32

1.3 C = -32

C = -24.6, F = -12.3

-24.6 degrees C = -12.3 degrees F. That’s rounded to the nearest tenth of a degree.

At what temperature the Fahrenheit scale reading equal to half of Celsius scale:

Basic conversion formula: F = 9/5C + 32
Our problem: 2C = F

Substitute F in the first formula with 2C from the second, which gives you:
2C = 9/5C + 32

Subtract 9/5C from both sides, which gives:
2C - 9/5C = 32

Simplify the left side (2C = 10/5C): 10/5C - 9/5C = 32, which equals 1/5C = 32

Multiply 5 to each side, which leaves C = 160

Kushagra Madhukar
3 years ago
As we know, the formula for interconversion between Fahrenheit and Celsius is 
F = 9/5 C + 32
 
(a) given, F = 2C
hence,
2C = 9/5 C + 32
or, C/5 = 32
or, C = 160o celsius which in Fahrenheit scale, F = 320
 
(b) given, F = C/2
hence, C/2 = 9/5 C + 32
or, – 2/5 C = 32
or, C = – 80o celsius; which in fahrenheit scale is, F = -40
 
Hope it helps.
Regards,
Kushagra